### v2

@@ 1 Introduction
Figure 1: Overview of our approach. Our key insight is to divide the vocabulary into context-independent and context-dependent tokens at each position within the pushdown automaton. We precompute and cache the context-independent tokens in an adaptive token mask cache, which is then retrieved at runtime. Other context-dependent tokens are checked on the fly. Additionally, we implement various optimizations to reduce the number of context-dependent tokens and enhance processing efficiency, ultimately accelerating runtime handling of these tokens.
Recent advancements in large language models (LLMs) have created new possibilities for complex applications such as code generation Chen et al. (2021); Wang et al. (2021), debugging Pearce et al. (2022); Mozannar et al. (2024), external tool invocation through function calling OpenAI (2024); LangChain (2024), and robotic control Liu et al. (2023). These applications bring great demand for LLM systems to perform structured generation and produce outputs that follow specific formats, such as JSON, SQL or other structures tailored to the task. The downstream applications can then organically consume the structured outputs to perform followup interactions with the system.
Constrained decoding Deutsch et al. (2019); Kuchnik et al. (2023) is a commonly adopted method for structured generation. At each decoding step, constrained decoding examines the vocabulary and filters out tokens that violate the specified structure by setting the probabilities of invalid tokens to zero. To support the rich structure formats arising in diverse applications, a flexible mechanism is needed to specify and check the constraints. Context-free grammar (CFG) Chomsky (1956); Poesia et al. (2022); Scholak et al. (2021) provides a general approach for defining structures through a set of rules. Each rule contains a sequence of characters or other rules, allowing recursive composition to represent complex structures. Compared to alternative formats such as regular expressions, CFGs offer greater flexibility by allowing recursive structures, making them suitable for describing common languages such as JSON, SQL, and domain-specific languages (DSLs).
@@ 2.1 LLM Constrained Generation
Figure 3: Up: A context-free grammar for arrays and strings that can be recursively composed. This CFG is converted into the pushdown automata in Figure 1. [^"\] denotes every character except " and \. Down: Two possible matching stacks for matching the string ["a to the CFG. Each stack represents a possible expansion of the rules in the CFG.
@@ 2.2 Context-free Grammar and Pushdown Automata
Pushdown automata (PDA) Schützenberger (1963); Evey (1963) are typically used to recognize languages generated by CFGs, as they employ a stack to manage nested structures. An example of PDA is shown in Figure 1, and its stacks are shown in detail in Figure 3. A PDA consists of multiple finite state automata (FSA), each representing a grammar rule, with the stack handling recursive rule expansions. The transitions in the FSA include two types: character edges, which accept specific characters, and rule reference edges, which allow recursive entry into other rules. To match a string, the PDA begins with the main rule, recursively expanding child rules by pushing rule-reference edges onto the stack; once a rule is fully matched, it pops the stack to return to the previous rule. The top of the stack holds the current node reached. If the grammar is ambiguous, meaning it allows multiple rule expansions for the same input string, the PDA can maintain multiple parallel stacks for each expansion path, ensuring flexibility. However, the unbounded stack length results in an infinite number of possible states, making it impractical to precompute token masks for all scenarios, thus posing challenges for efficient constrained decoding.
@@ 3.1 Adaptive Token Mask Cache
Thus, in both accept-heavy and reject-heavy cases, the adaptive storage format only requires storing a small subset of tokens, significantly reducing memory usage. For Llama-3.1 model and JSON grammar, this adaptive storage method can effectively reduce the total memory usage to 0.2% (from 160 MB to 0.46 MB).
@@ 3.2 Context Expansion
Figure 6: The context expansion. Each rule obtains a set of expanded suffices, representing the set of strings that must be matched after completing this rule. If a context-dependent token cannot match any of these suffices, it is rejected.
@@
When validating a context-dependent token on a pushdown automaton, the matching process will reach the end of the current rule and return to parent rules to continue checking. This means a prefix of the token can be matched by the current rule, but whether the rest part can be matched remains to be determined by parent rules. However, we observed that the set of possible parent rules for each rule is limited, and the set of strings that can continue be matched after returning to parent rules is often constrained. Based on this observation, context expansion precomputes the possible suffix strings for each rule when returning to parent rules, called the expanded suffix. If a context-dependent token cannot match any string in the expanded suffix after finishing the current rule, it is marked as invalid. This filtering process effectively reduces the number of context-dependent tokens by eliminating those that would fail in higher-level rule contexts. Applied to the Llama-3.1 model and JSON grammar, this technique reduces context-dependent tokens by 90% (from 1,134 to 120), substantially improving the efficiency of generating token masks at runtime.
@@ 3.3 Persistent Execution Stack
There are many LLM applications that involve rolling back the output to a previous token. For instance, the jump-forward decoding requires retokenization, which involves rolling back some tokens in the context and then inserting new tokens. To ensure structured generation can continue after rolling back tokens, we can roll back the automaton state simultaneously with the output token rollback. There are also many LLM applications that requires LLMs generate in a tree structure, such as in Tree-of-thought Yao et al. (2024), SGLang Zheng et al. (2024), and the speculative model in the speculative decoding algorithm SpecInfer Miao et al. (2024). We can maintain the automata state for every branch of the output tree, and when the output branches, we can quickly split the automaton state, maintaining separate matching states for each output branch. This branching is fast because we only need to maintain the stack top pointer on the tree for every branch. Therefore, the persistent execution stack enables us to ensure efficient structured generation for all these applications.
@@ 4 Evaluation
  • Can XGrammar be deployed across a broader range of platforms? (§11)
@@
4.1 Grammar Engine Efficiency
This subsection evaluates the grammar engine performance. We evaluate our method and baselines on Llama-3.1-8B-Instruct, a popular model with the ability to follow human instructions. We first evaluate the performance of JSON grammar. We apply the standard context-free grammar of JSON adopted from ECMA-404 Ecma International (2013) as a context-free grammar without additional constraints. We also evaluate the JSON schema, where we leverage the additional schema constraints from the dataset. We utilize the JSON-mode-eval dataset NousResearch (2024) for the prompts. We run the evaluation AMD Ryzen 9 7950X CPU and NVIDIA RTX 4090 GPU. For baseline comparisons, we compare three two popular implementations of structured generation engine, Outlines Willard & Louf (2023)(v1.0) and the builtin grammar engine of llama.cpp Gerganov (2023) (b3998).
The results are shown in Figure 9. XGrammar can achieve up to 3x speedup in the setting of JSON schema, and more than 100x speedup in the case of JSON grammar. The context-free grammar of JSON contains more complicated rules compared to the JSON schema (which is more constrained), as it can contain recursive lists and dictionaries, making it harder for grammar engines to execute it efficiently. In both cases, XGrammar can generate each token mask at in less than 40us, making it ideal for low-latency LLM inference.
4.2 End-to-End LLM Engine Evaluation
This section evaluates XGrammar under LLM serving setting. We integrate XGrammar into an end-to-end LLM inference framework and compare its efficiency with other LLM serving frameworks. We also integrate XGrammar into SGLang, an LLM serving engine to demonstrate its ability to efficiently integrate into LLM frameworks. We measure the average time to the first token (TTFT), which is primarily affected by preprocessing the constraint, and the average time per output token (TPOT), which is primarily affected by applying the constraint to each output token. We compare the efficiency with other LLM engines that support structured generation, including vLLM Kwon et al. (2023b)(v0.6.3) integrated with Outlines, and llama.cpp with its builtin grammar engine. We conduct the evaluations on Llama-3.1-8B-Instruct under JSON grammar and JSON schema. We turn on the grammar cache for all engines to enable caching of the preprocessed grammars. The hardware used for the tests is AMD EPYC 7R13 CPU and NVIDIA H100 GPU. We evaluate multiple batch sizes settings in LLM inference tasks.
Table 1: Comparison of the Llama3.1 TPOT (ms) for the XGrammar engine, with and without grammar constraint enabled.
Task Batch Size Constraint Constraint
Off On
@@ 4.2 End-to-End LLM Engine Evaluation
The experiment results are shown in Figure 10. XGrammar achieves the best TTFT and TPOT among all baselines for both CFG and JSON Schema. The computation of vLLM and llama.cpp is hindered by their grammar engines’ longer preprocessing and per-token processing times. The decrease in TPOT speed in vLLM becomes particularly noticeable with larger batch sizes. This is because a larger batch size leads to higher throughput, putting greater pressure on grammar processing. Overall XGrammar engine can bring up to 80x speed output token rate compared to existing solutions. This speedup comes from the performance optimizations bought by XGrammar. We also studied the overhead of grammar processing in Table 1. The grammar process incurs nearly zero overhead in the TPOT, thanks to the token mask generation efficiency and grammar GPU overlap.
4.3 Cross-platform Deployment
This section explores bringing XGrammar to a wide variety of platforms. We leverage Emscripten  Zakai (2011) to compile XGrammar into WebAssembly  Haas et al. (2017) and build a JavaScript binding. This approach enables XGrammar to run in client-side browsers on portable devices like laptops and mobile phones. We further integrate the web-binding with the in-browser LLM inference framework WebLLM  MLC team (2023b) to enable structured generation. We evaluate the end-to-end performance with the JSON-mode-eval dataset, using 4-bit quantized models Llama-3.1-8B-Instruct Dubey et al. (2024b) on a Macbook Pro M3 Max (macOS 14.5) with Google Chrome, and Qwen-2.5-0.5B-Instruct Yang et al. (2024) on an iPhone 14 Pro Max (iOS 18) with Safari.
@@ 5 Related Work
Outlines Willard & Louf (2023) and SynCode Ugare et al. (2024) utilize a lexer and parser to handle output and generate the token mask. However, these approaches can suffer from boundary mismatch problem Koo et al. (2024) problem. Synchromesh Poesia et al. (2022) and llama.cpp Gerganov (2023) check all tokens during runtime, leading to significant overhead. lm-format-enforcer Gat (2024) design optimizations for regular expressions that cannot be easily extended to context-free grammar. XGrammar brings a series of system optimizations to reduce the runtime check via context-independent caching to reduce per token generation. It also enables co-optimizations to enable end-to-end LLM inference speedup in structured generation settings.
@@
LLM serving engines MLC team (2023a); Zheng et al. (2024); Kwon et al. (2023a); hiworldwzj et al. (2024) employ various techniques to support efficient LLM generation for multiple concurrent users, including engine-level techniques such as continuous batching Yu et al. (2022) for dyanmic request scheduling, and low-level KV cache technique PagedKVCache Kwon et al. (2023a) for efficient memory management. These LLM serving engines can leverage XGrammar for efficient, structured generation on top of their existing LLM inference techniques.
@@ Acknowledgements
We thank (alphabetically) the DeepSeek team, SGLang team, TensorRT-LLM team, vLLM team, and WebLLM team for their helpful feedback and discussions. We also thank Weihua Du, Haoran Peng, Xinyu Yang, Zihao Ye, Zhihao Zhang, and Ligeng Zhu for their insightful discussion and feedback.

### v3

@@ 1 Introduction
Figure 1: Overview of our approach. XGrammar first uses a pushdown automaton to parse the prior LLM output, flexibly supporting diverse grammars and producing the matching stack states. It then uses the stack top to index into the adaptive token mask cache—our key optimization—to retrieve a partial mask. Most of the partial mask consists of context-independent tokens and is determined during preprocessing. A small portion, however, is context-dependent and resolved at runtime. This yields the complete token mask, thus enabling efficient constraint decoding.
Recent advancements in large language models (LLMs) have created new possibilities for complex applications such as code generation Chen et al. (2021); Wang et al. (2021), debugging Pearce et al. (2022); Mozannar et al. (2024), external tool invocation through function calling OpenAI (2024); LangChain (2024), and robotic control Liu et al. (2023). These applications bring great demand for the structured generation problem, which requires the output of LLMs conforms to specific formats or grammars, such as JSON, SQL or other formats tailored to the task. The downstream applications can then organically consume the structured outputs to perform followup interactions with the system.
Constrained decoding Deutsch et al. (2019); Kuchnik et al. (2023) is a commonly adopted method for structured generation. It guarantees the output of LLMs adheres to the specified structure through only allowing tokens that conform to the structure to be generated at each decoding step. At each step, constrained decoding first scans the entire vocabulary to identify invalid tokens and sets their probabilities to zero, thereby preventing them from being generated. To support the rich structure formats arising in diverse applications, a flexible mechanism is needed to specify and check the constraints. Context-free grammar (CFG) Chomsky (1956); Poesia et al. (2022); Scholak et al. (2021) provides a general approach for defining structures through a set of rules. Each rule contains a sequence of characters or other rules, allowing recursive composition to represent complex structures. Compared to alternative formats such as regular expressions, CFGs offer greater flexibility by allowing recursive structures, making them suitable for describing common languages such as JSON, SQL, and domain-specific languages (DSLs).
@@ 2.1 LLM Constrained Generation
Figure 3: Up: A context-free grammar for arrays and strings that can be recursively composed. This CFG is converted into the pushdown automata in Figure 1. [^"\] denotes every character except " and \. Down: Two possible matching stacks for matching the string ["a to the CFG. Each stack represents a possible expansion of the rules in the CFG. The edges and nodes in the stack correspond to the transitions and states in the PDA in Figure 1.
@@ 2.2 Context-free Grammar and Pushdown Automata
Pushdown automata (PDA) Schützenberger (1963); Evey (1963) are typically used to recognize languages generated by CFGs, as they employ a stack to manage nested structures. In this paper, we use a definition of PDA that is equivalent to the original one, but more conducive to explaining the algorithm. An example of PDA is shown in Figure 1, and its stacks are shown in detail in Figure 3. A PDA consists of multiple finite state automata (FSA), each representing a grammar rule, with the stack handling recursive rule expansions. The transitions in the FSA include two types: character edges, which accept specific characters, and rule reference edges, which allow recursive entry into other rules. A formal definition of the PDA is provided in Appendix A. To match a string, the PDA begins with the main rule, recursively expanding child rules by pushing rule-reference edges onto the stack; once a rule is fully matched, it pops the stack to return to the previous rule. The top of the stack holds the current node reached. If the grammar is non-deterministic, meaning there can be multiple possible transitions in the PDA for the same input character, the PDA can maintain multiple parallel stacks for each path, ensuring flexibility. However, the unbounded stack length results in an infinite number of possible states, making it impractical to precompute token masks for all scenarios, thus posing challenges for efficient constrained decoding.
@@ 3.1 Adaptive Token Mask Cache
Thus, in both accept-heavy and reject-heavy cases, the adaptive storage format only requires storing a small subset of tokens, significantly reducing memory usage. In practice, we will enumerate the three storage types, calculate their respective costs, and choose the storage type with the smallest size. For Llama-3.1 model and JSON grammar, this adaptive storage method can effectively reduce the total memory usage to 0.2% (from 160 MB to 0.46 MB).
@@ 3.2 Context Expansion
Figure 6: The context expansion. Each rule obtains a set of expanded suffices, representing the set of strings that must be matched after completing this rule. The remaining unmatched part of the context-dependent tokens should either be a prefix of the expanded suffix or start with the expanded suffix. Otherwise, they are rejected.
@@
As described in the last section, a token is context-dependent when we reach the end of the current rule during matching, but there is still a remaining part of the token that requires further checking by returning to the parent rules. However, through analysis of the grammar, we can observe that in a large portion of the cases, the remaining part of a token is invalid. This is because, for many rules, when they reach their end, there are only a limited number of positions they can return to within their parent rules, and the set of strings that can be further matched from those positions is also limited.
Based on this observation, context expansion precomputes, for each rule, the set of strings that can be accepted after returning to the parent rules, called the expanded suffix. The remaining unmatched part of the context-dependent tokens should either be a prefix of the expanded suffix or start with the expanded suffix. Otherwise, they are rejected. This filtering process effectively reduces the number of context-dependent tokens by eliminating those that would fail in higher-level rule contexts. Applied to the Llama-3.1 model and JSON grammar, this technique reduces context-dependent tokens by 90% (from 1,134 to 120).
@@ 3.3 Persistent Execution Stack
There are many LLM applications that involve rolling back the output to a previous token. For instance, the jump-forward decoding Yin et al. (2024) requires retokenization, which involves rolling back some tokens in the context and then inserting new tokens. To ensure structured generation can continue after rolling back tokens, we can roll back the automaton state simultaneously with the output token rollback. There are also many LLM applications that requires LLMs generate in a tree structure, such as in Tree-of-thought Yao et al. (2024), SGLang Zheng et al. (2024), and the speculative model in the speculative decoding algorithm SpecInfer Miao et al. (2024). We can maintain the automata state for every branch of the output tree, and when the output branches, we can quickly split the automaton state, maintaining separate matching states for each output branch. This branching is fast because we only need to maintain the stack top pointer on the tree for every branch. Therefore, the persistent execution stack enables us to ensure efficient structured generation for all these applications.
@@ 4 Evaluation
  • How effective is each optimization technique introduced in XGrammar? (§4.3)
  • How does XGrammar effect downstream structured generation tasks? (§4.4)
4.1 Mask generation efficiency
This section evaluates the efficiency of mask generation to measure the overhead introduced by constraint decoding. We first assess regex-based methods using JSON schemas, which can be converted into regex. To test more complex cases beyond regex capabilities, we evaluate context-free grammars, including unconstrained JSON (from ECMA-404 Ecma International (2013)), XML (based on the XML 1.0 standard Bray et al. (2008)), and a Python DSL (adapted from the Python Grammar Specification Python Software Foundation (2024)). Unconstrained JSON cannot be handled by regex-based methods due to its support for arbitrarily nested lists and objects. The Python DSL covers basic control flow (if, for, while) and data types (str, int, float, bool) but ignores indentation. For JSON schema and unconstrained JSON, we use the JSON-mode-eval dataset NousResearch (2024), and for XML and Python, we use a synthetic dataset. For baselines, we choose three popular constrained generation libraries: Outlines Willard & Louf (2023) (v1.0), the grammar engine in llama.cpp Gerganov (2023) (b3998), and lm-format-enforcer Gat et al. (2025) (v0.10.9, a regex-based method that does not support CFG). All methods are evaluated on Llama-3.1-8B-Instruct using an AMD Ryzen 9 7950X CPU and an NVIDIA RTX 4090 GPU.
The results are shown in Figure 9. XGrammar consistently achieves the lowest latency across all tasks, with under 40 µs per token for JSON Schema and CFG (JSON), and under 200 µs for XML and Python DSL. It delivers up to 3x speedup on JSON Schema and over 100x on CFG, compared to the best baseline in each case.
Figure 9: Per token mask generation latency. XGrammar consistently outperforms existing constrained decoding libraries.
4.2 End-to-End LLM Engine Evaluation
@@
Table 1: End-to-end structured generation efficiency across different models, measured in time per output token (ms) on the JSON Schema task and the Llama-3.1 8B model. XGrammar demonstrates superior performance across different models.
Model SGLang SGLang
+ Outlines + XGrammar
Llama-3.1 8B 44.2 6.8
DeepSeek-V2-Lite 15.8 4.8
16B MOE
This section evaluates XGrammar in LLM serving scenarios. We co-design XGrammar with the serving engines using the overlapping technique introduced in §3.5. We integrate it into widely used end-to-end LLM serving engines, including the C++-based MLC-LLM MLC team (2023a) and the Python-based SGLang Zheng et al. (2024), showcasing XGrammar’s adaptability and efficiency across different deployment environments.
We first compare the efficiency of several LLM engines that support structured generation, including vLLM Kwon et al. (2023b) (v0.6.3) with Outlines and llama.cpp with its built-in grammar engine. Efficiency is measured by the average time per output token (TPOT), which reflects the overhead of applying constraints during token generation. The evaluations are conducted using Llama-3.1-8B-Instruct under both JSON schema and CFG (unconstrained JSON). All tests are run in an online serving setting with fixed batch sizes, on hardware with an AMD EPYC 7R13 CPU and an NVIDIA H100 GPU. To ensure a fair comparison, we set a fixed maximum output length. On average, the input has 139 tokens, and the generated output has 53 tokens.
The experiment results are shown in Figure 10. XGrammar achieves the best TPOT among all baselines for both JSON Schema and CFG. The computation of vLLM and llama.cpp is hindered by their grammar engines’ longer preprocessing and per-token processing time. The decrease in TPOT speed in vLLM becomes particularly noticeable with larger batch sizes. This is because a larger batch size leads to higher throughput, putting greater pressure on grammar processing on the CPU side. Overall the XGrammar-based structured generation solutions can bring up to 80x output token rate compared to existing solutions. This proves the effectiveness of XGrammar’s system optimizations and its co-design with the serving engine.
We also study the effectiveness of XGrammar across different models in end-to-end serving scenarios. As the result shown in Table 1, for various models, SGLang integrated with XGrammar consistently outperforms its integration with Outlines, demonstrating the robustness of XGrammar’s effectiveness across model architectures. This experiment also provides direct evidence that XGrammar outperforms other constrained decoding library when running on the same serving engine.
Additionally, we examine the overhead of XGrammar’s constrained decoding in end-to-end scenarios by measuring performance on the MLC-LLM engine with and without XGrammar. As shown in Table 2, enabling XGrammar enhances output quality with nearly zero overhead in TPOT. This is attributed to efficient mask generation and the overlapping of grammar processing with GPU execution.
Table 2: The impact on performance of enabling and disabling XGrammar tested on the MLC-LLM Engine and the Llama-3.1 8B model. TPOT (ms) is reported. XGrammar introduces minimal overhead to the serving engine with better generation quality.
Task Batch TPOT w/o TPOT w/
Size XGrammar XGrammar
@@
4.3 Ablation Study of Optimization Techniques
Table 3: Ablation study of optimization techniques in XGrammar.
Optimization Per-token latency (ms)
PDA Baseline 65.776
+ Node merging 38.280 (1.7×↓)
+ Adaptive token mask cache 0.154 (248.6×↓)
+ Rule inlining 0.035 (4.4×↓)
+ Context Expansion 0.018 (1.9×↓)
In this section, we investigate the impact of various optimizations introduced in xgrammar on mask generation performance, to better illustrate our design decisions. We begin by implementing a baseline using a pushdown automaton parser without any optimizations, where each token mask is generated by checking the entire vocabulary to determine whether parsing can proceed. Building on this baseline, we progressively add the optimizations described in this paper, namely node merging, the adaptive token mask cache, rule inlining, and context expansion. For each configuration, we measure the average mask generation time on the CFG (unconstrained JSON) task with the Llama-3.1 8B model and the json-mode-eval dataset. As shown in Table 3, the results show that the adaptive token mask cache has the greatest impact in terms of speedup, while other techniques, including node merging, rule inlining, and context expansion, also yield noticeable improvements.
4.4 Impact of XGrammar on Structured Generation Tasks
XGrammar can improve the generation quality of LLMs by ensuring that the output strictly adheres to the given format. We evaluate the impact of XGrammar on two strcutured generation tasks: function calling (i.e., JSON generation guided by a JSON schema) and XML code generation. For function calling, we use the json-mode-eval dataset, while for XML code generation, we rely on a synthetic dataset. We measure the syntactic correctness of the generated function calling outputs and XML code using LLaMA-3.1 8B. As shown in Table 4, XGrammar significantly improves generation accuracy. We observe that without XGrammar, the model often includes additional explanations alongside the intended code output, or the generated JSON contains an unexpected type. This makes the output unsuitable for direct use by downstream applications. XGrammar avoids this issue by enforcing grammar constraints.
Table 4: Impact of XGrammar on structured generation tasks. XGrammar ensures 100% syntactic correctness of the generated outputs.
Task Accuracy w/o Accuracy w/
XGrammar XGrammar
Function calling 62% 100%
XML code generation 80% 100%
@@ 5 Related Work
There has also been some previous work focusing on constrained decoding. Some methods use regex to represent syntax, such as lm-format-enforcer Gat (2024), but they cannot handle more complex CFGs. Synchromesh Poesia et al. (2022) and llama.cpp Gerganov (2023) utilizes LR parser and PDA respectively to handle CFG and generate the mask, but requires checking the entire vocabulary at runtime, which incurs large overhead. Outlines Willard & Louf (2023) handles the grammar through a lexer and parser, and uses caching to accelerate mask generation, but it only considers the most recent lexer token, which may lead to incorrect judgments when LLM tokens span across multiple lexer tokens (Appendix A in Koo et al. (2024)). Syncode Ugare et al. (2024) uses a lexer and parser with a cache spanning multiple lexer tokens, but requires all tokens to be processed offline, leading to substantial preprocessing overhead. In this work, XGrammar leverages a PDA to support CFGs and introduces an adaptive token mask cache to adaptively handle tokens during preprocessing and runtime, combining other system optimizations to achieve minimal overhead in both stages.
@@
LLM serving engines MLC team (2023a); Zheng et al. (2024); Kwon et al. (2023a); hiworldwzj et al. (2024) employ various techniques to support efficient LLM generation for multiple concurrent users, including engine-level techniques such as continuous batching Yu et al. (2022) for dyanmic request scheduling, and low-level KV cache technique PagedKVCache Kwon et al. (2023a) for efficient memory management. Also, AICI Moskal et al. (2024) proposes a CPU-GPU parallel computation paradigm to accelerate structured generation. These LLM serving engines can leverage XGrammar for efficient, structured generation on top of their existing LLM inference techniques.
@@ Acknowledgements
This work was supported in part by NSF award CNS-2211882, and gifts from OctoAI, Qualcomm, and CMU opensource software fellowships. We are grateful to Sasa Misailovic for his guidance as the shepherd of this paper. We also appreciate the valuable feedback and discussions from the DeepSeek, SGLang, TensorRT-LLM, vLLM, and WebLLM teams (listed alphabetically). Additionally, we thank Weihua Du, Haoran Peng, Xinyu Yang, Zihao Ye, Jieyu Zhang, Zhihao Zhang, and Ligeng Zhu for their insightful input and thoughtful conversations.
Appendix A Formal Definition of the PDA Variant
In our paper, we define a variant of the pushdown automaton (PDA) that is equivalent to the original definition, but is designed to facilitate the description of the parsing algorithm and the construction of the token mask cache, since the keys of the token mask cache are precisely the states in this PDA. It is defined as the tuple
$P=\bigl{(}R,\Sigma,\{A_{r}\}_{r\in R},q_{\text{main}},\delta\bigr{)},$
where:
  • $R$ is a finite set of grammar rules.
  • $\Sigma$ is a finite input alphabet.
  • For each rule $r\in R$, the corresponding finite state automaton is given by
    $A_{r}=\bigl{(}Q_{r},\Sigma\cup R,q^{\text{start}}_{r},F_{r},\delta_{r}\bigr{)},$
    where $Q_{r}$ is a finite set of states, $q^{\text{start}}_{r}\in Q_{r}$ is the start state, $F_{r}\subseteq Q_{r}$ is the set of accepting states, and $\delta_{r}$ is the transition function defined over $Q_{r}$. The transition labels in $A_{r}$ are drawn from the alphabet $\Sigma\cup R$, which includes both input characters and rule references.
  • $q_{\text{main}}$ is the start state corresponding to the main rule.
  • $\delta$ is the global transition function that governs the operation of the PDA by handling two kinds of transitions:
[…]
    • Rule reference transitions: When in a state $q\in Q_{r}$, a transition labeled by a rule reference to $s\in R$ allows the PDA to push the current return information onto the stack and jump to the start state $q^{\text{start}}_{s}$ of $A_{s}$.
The parsing state is represented by a set of pairs $\{(s_{i},q_{i})\}$, where $s_{i}$ denotes the content of the stack (encoding return information), $q_{i}$ is the current state (with $q_{i}\in Q_{r}$ for some $r\in R$). There can be multiple such pairs because the pushdown automaton can contain non-deterministic transitions, which means there could be multiple possible parsing stacks and states. In the main body of the paper, for simplicity, we place the current state $q_{i}$ at the top of the stack. Thus, the parsing state is represented as a set of parsing stacks.
We now describe a formal transformation that converts the above variant PDA definition into a standard PDA. To obtain a standard PDA, we construct a new pushdown automaton
$P^{\prime}=\bigl{(}Q,\Sigma,\Gamma,\delta^{\prime},q_{0},F\bigr{)}$
as follows. The state set $Q$ is defined as the union of all states from the FSAs:
$Q=\bigcup_{r\in R}Q_{r},$
and we set the initial state to be $q_{0}=q_{\text{main}}$. The stack alphabet $\Gamma$ is chosen to record return information; we define
$\Gamma=\{\,(r,q)\mid r\in R,\;q\in Q_{r}\,\},$
so that each symbol $(r,q)$ encodes the context of a rule call, namely the originating rule and the return state.
The transition function $\delta^{\prime}$ of the standard PDA is then defined to simulate the behavior of the variant PDA. For each character transition in some $\delta_{r}$, if
[…]
we include in $\delta^{\prime}$ the transition
$\delta^{\prime}(q,a,\gamma)\ni(q^{\prime},\gamma)\quad\text{for all }\gamma\in\Gamma.$
For each rule reference transition in $\delta_{r}$, if
[…]
we simulate the recursive call by defining an $\epsilon$-transition that pushes the return information and transfers control to the called rule. Formally, we set
$\delta^{\prime}(q,\epsilon,\gamma)\ni\Bigl{(}q^{\text{start}}_{s},\,(r,q^{\prime})\cdot\gamma\Bigr{)}\quad\text{for all }\gamma\in\Gamma,$
where $q$ belongs to the automaton $A_{r}$, and $(r,q^{\prime})\cdot\gamma$ denotes the stack obtained by pushing $(r,q^{\prime})$ onto $\gamma$. Finally, when an automaton $A_{r}$ reaches an accepting state $q\in F_{r}$, the standard PDA simulates the return from a recursive call by popping the top of the stack. That is, if the current stack has the form $(r^{\prime},q^{\prime})\cdot\gamma$, we define
$\delta^{\prime}(q,\epsilon,(r^{\prime},q^{\prime}))\ni(q^{\prime},\gamma).$
The set of accepting states $F$ is defined as those states in $\bigcup_{r\in R}F_{r}$ that are reached with an empty stack.
This construction shows that character transitions in the variant PDA directly correspond to state transitions without stack operations in $P^{\prime}$, while rule reference transitions correspond to stack push and jump operations. Similarly, completing the match of an FSA and returning to the parent rule is achieved by a stack pop and a transition to the stored return state. In this way, the standard PDA $P^{\prime}$ exactly simulates the recursive behavior of our variant PDA.
Appendix B Synergy between XGrammar and Jump-forward Decoding
Jump-forward decoding is a technique designed to accelerate structured generation. When, based on the current input, the following output can be deterministically inferred from the grammar, it bypasses LLM decoding and sampling by directly tokenizing and appending the output to the context. This improves the end-to-end efficiency of LLM generation. This technique is orthogonal to the constrained decoding adopted by XGrammar. XGrammar further supports jump-forward decoding and demonstrates that the two can be effectively combined to further improve efficiency.
We evaluate performance using the Llama-3.1-8B-Instruct model served through the SGLang engine, running on a machine with an AMD Ryzen 9 7950X CPU and an NVIDIA RTX 4090 GPU. All experiments are conducted with a batch size of 1. We compare our method, XGrammar, against the existing decoding backend Outlines, testing both with and without jump-forward decoding enabled. We measure time per output token as the evaluation metric.
As shown in Figure 11, XGrammar consistently outperforms Outlines, and achieves the best efficiency when combined with jump-forward decoding—demonstrating its ability to better leverage structural constraints for faster generation.
Figure 11: The performance comparison with and without jump-forward decoding on the JSON schema task and the SGLang engine. XGrammar combined with jump-forward can achieve optimal TPOT.
Appendix C Cross-platform Deployment of XGrammar
We study bringing XGrammar to a wide variety of platforms. We leverage Emscripten  Zakai (2011) to compile XGrammar into WebAssembly  Haas et al. (2017) and build a JavaScript binding. This approach enables XGrammar to run in client-side browsers on portable devices like laptops and mobile phones. We further integrate the web-binding with the in-browser LLM inference framework WebLLM  MLC team (2023b) to enable structured generation.
We evaluate the end-to-end performance with the JSON-mode-eval dataset, using 4-bit quantized models Llama-3.1-8B-Instruct Dubey et al. (2024b) on a MacBook Pro M3 Max (MacOS 14.5) with Google Chrome, and Qwen-2.5-0.5B-Instruct Yang et al. (2024) on an iPhone 14 Pro Max (iOS 18) with Safari.